### v1

@@ Abstract
Robust adversarial reinforcement learning has emerged as an effective paradigm for training agents to handle uncertain disturbance in real environments, with critical applications in sequential decision-making domains such as autonomous driving and robotic control. Within this paradigm, agent training is typically formulated as a zero-sum Markov game between a protagonist and an adversary to enhance policy robustness. However, the trainable nature of the adversary inevitably induces non-stationarity in the learning dynamics, leading to exacerbated training instability and convergence difficulties, particularly in high-dimensional complex environments. In this paper, we propose a novel approach, Uncertainty-Aware Critic Ensemble for robust adversarial Reinforcement learning (UACER), which consists of two strategies: 1) Diversified critic ensemble: a diverse set of $K$ critic networks is exploited in parallel to stabilize Q-value estimation rather than conventional single-critic architectures for both variance reduction and robustness enhancement. 2) Time-varying Decay Uncertainty (TDU) mechanism: advancing beyond simple linear combinations, we develop a variance-derived Q-value aggregation strategy that explicitly incorporates epistemic uncertainty to dynamically regulate the exploration-exploitation trade-off while simultaneously stabilizing the training process. Comprehensive experiments across several MuJoCo control problems validate the superior effectiveness of UACER, outperforming state-of-the-art methods in terms of overall performance, stability, and efficiency.
keywords:
[…]
1 Introduction
@@
Deep reinforcement learning (DRL) has established itself as a transformative paradigm for sequential decision-making, demonstrating remarkable success beyond gaming domains. Its applications now span real-world scenarios including precision robotic control (Huang et al., 2022; Cui et al., 2022; Valencia et al., 2024; Tang et al., 2025), adaptive healthcare systems (Yu et al., 2021a; Wang et al., 2022; Luo et al., 2024), and autonomous vehicle navigation (Feng et al., 2023; Sun et al., 2023; Zhou et al., 2024). However, a critical gap persists: most DRL policies are trained in idealized simulated environments that fail to capture the complexity of physical world deployment. This issue becomes particularly pronounced in certain high-dimensional complex tasks such as robotic control. Even minor discrepancies in physical parameters, such as friction, mass distribution, or object appearance, can lead to significant performance deterioration when policies are deployed to real-world systems. More specifically, these policies frequently overfit to simulation-specific dynamics, resulting in catastrophic performance declines when confronted with inevitable real-world disturbances (Gleave et al., 2020) or distribution shifts (Chen et al., 2021a; Zhang et al., 2024). Addressing this vulnerability, recent research (Pinto et al., 2017; Zhang et al., 2020a; Reddi et al., 2024) has prioritized the development of RL frameworks capable of counteracting adversarial disturbances and environmental variations.
@@
Robust adversarial reinforcement learning (Pinto et al., 2017; Sheng et al., 2022; Reddi et al., 2024) is a prominent class of methods in this research domain, specifically designed to enhance the stability and generalization capabilities of RL agents operating in uncertain or adversarially perturbed environments. The framework reformulates conventional RL as a two-player zero-sum Markov game (Littman, 1994; Perolat et al., 2015), comprising a protagonist agent that optimizes policy parameters to maximize cumulative returns, and an adversary agent that systematically generates worst-case adversarial perturbations to minimize these returns. While this minimax formulation theoretically enables agents to develop resilience against worst-case perturbations, its practical deployment poses significant methodological challenges (Zhang et al., 2020b): (i) Non-Stationary Training Dynamics: The adversary’s evolving policy induces continuous perturbations, dynamically altering the protagonist’s observed state distribution and reward signals. This interaction creates a highly non-stationary learning environment where the protagonist’s optimization target shifts iteratively. As a result, policy updates exhibit high variance, frequently destabilizing the training process. (ii) Convergence Impediments: The cyclical adversarial training regime forces the protagonist to adapt to an ever-changing adversary policy before its own policy converges. Without a fixed optimization target, the protagonist remains trapped in a persistent reactive adaptation loop, leading to significantly prolonged convergence times due to continual policy adjustments, or even completed convergence failure in cases of extreme non-stationary conditions.
@@
To address the aforementioned stability and convergence problems, we propose Uncertainty-Aware Critic Ensemble robust adversarial Reinforcement learning (UACER), a simple yet effective general framework that advances conventional robust adversarial RL by systematically integrating ensemble methodology with uncertainty quantification. This integration significantly improves both training stability and convergence speed in adversarial environments. The technical foundation of UACER addresses a fundamental limitation of single-critic architectures, which are particularly vulnerable to adversarial perturbations and distributional outliers, leading to unstable Q-value estimation (An et al., 2021a; Peer et al., 2021). To overcome this limitation, UACER employs a parallel ensemble of critic networks that provides three key benefits: Firstly, aggregated Q-value reduces the variance in the Q-function estimate and policy gradient updates so as to enhance training stability. Secondly, it effectively mitigates training oscillations and increases resilience against adversarial disturbances through diversified value function approximations. Thirdly, the increased stability and robustness facilitate accelerated policy convergence during iterative adversarial interactions. For the final Q-value determination session, we present a Time-varying Decay Uncertainty (TDU) Q-value aggregation function for combining outputs across multiple critic networks. The TDU mechanism is designed to: (i) promote extensive exploration during initial training phases through higher uncertainty tolerance, and (ii) gradually shift towards more conservative value estimation in later stages to ensure robust convergence. This way, we can enable effective balance between exploration and exploitation together with maintaining resilience against adversarial perturbations throughout the learning process. Our contributions can be summarized as follows:
@@
    We develop a novel uncertainty-aware critic ensemble method for robust adversarial reinforcement learning (called UACER), which achieves more stable and accurate Q-value estimation through employing multiple base critic predictors.
@@
    We design a TDU function, a principled decision strategy to combine Q-value predictions from multiple critics. This mechanism dynamically balances exploration and exploitation while maintaining stable policy updates throughout the training process.
@@
    Quantitative experimental results on several complex MuJoCo problems demonstrate that UACER performs consistently better than state-of-the-art methods, including QARL and related baselines, in terms of overall performance, stability, and efficiency.
2 Related Work
Robust Adversarial Reinforcement Learning (RARL).
@@ 2 Related Work
The concept of adversarial training for robust RL was firstly established by RARL (Pinto et al., 2017), which formulated the problem as a zero-sum Markov game between a protagonist agent and an adversarial opponent. This paradigm demonstrated that exposing agents to strategically generated perturbations during training could significantly enhance their robustness against real-world distributional shifts. Building upon this framework, subsequent research has developed specialized improvements targeting several key limitations. For instance, RARARL (Pan et al., 2019) incorporated explicit risk modeling to reduce catastrophic failures, while NR-MDP (Tessler et al., 2019) expanded the adversary’s action space to generate more diverse and challenging disturbances. Further optimization advances introduced subsequently by MixedNE-LD (Kamalaruban et al., 2020) and Cen et al. (2021) employed stochastic gradient Langevin dynamics and extragradient methods, respectively, to improve the robustness of adversarial training. More recently, A2P (Liu et al., 2024) proposed dynamically adjusting the intensity of action-space perturbations based on the agent’s current relative performance, enabling more adaptive adversarial training. While these approaches have progressively addressed specific robustness challenges, they remain limited in handling the fundamental stability and convergence inherent in adversarial training. Our work advances this line of research by developing a unified framework that simultaneously improves training stability and convergence properties, while preserving and further enhancing the robustness benefits of adversarial training paradigms.
Training Stability and Convergence Improvements for RARL.
@@
While RARL has markedly advanced the robustness of RL policies, its adversarial formulation introduces fundamental challenges in training stability and convergence. These limitations are not confined to complex settings but are in inherent to the framework itself, as rigorously demonstrated by Zhang et al. (2020b), who showed their persistence even in simple linear quadratic regulation problems. In response, recent research has primarily evolved along three directions: (i) Constrained optimization approaches, where DI-CARL (Zhai et al., 2022) established stability guarantees through dissipation inequation constraint and $L_{2}$ gain performance in adversary RL; Yu et al. (2021b) proposed an accelerated convergence method for linear quadratic games by employing Bregman divergence to effectively capture and adapt to the global structure of the constraint set. (ii) Curriculum learning methods enhance stability and robustness by progressively exposing the agent to escalating adversarial difficulty, notably through CAT (Sheng et al., 2022), which employed a structured curriculum to systematically increases task complexity, thereby ensuring stable and efficient training process. (iii) hybrid frameworks such as QARL (Reddi et al., 2024) that combine entropy regularization with quantal response equilibrium and use a temperature parameter to systematically manage adversary rationality in a curriculum learning manner, thereby accelerating convergence while easing the complexity of the saddle point optimization problem. In contrast, our work adopts an ensemble-based perspective that provides inherent stability to adversarial RL architecture through diversified Q-value estimation. This approach maintains compatibility with existing optimization strategies and further enhances the stability and convergence properties.
Q-Function Ensemble in Reinforcement Learning.
@@
The integration of multiple Q-value estimators, known as Q-function ensemble, has become a widely adopted approach to enhance the stability and robustness of value-based deep RL. Early ensemble methods in the DQN family primarily address overestimation bias through averaged or pessimistic selection strategies, such as the simple averaging in Averaged DQN (Anschel et al., 2017) and the min-Q operator used in TD3 (Fujimoto et al., 2018), SAC (Haarnoja et al., 2018), and Maxmin Q-learning (Lan et al., 2020). This conservative estimation approach was further extended by methods like MSG (Ghasemipour et al., 2022) and EDAC (An et al., 2021b), which employed lower confidence bounds to systematically reduce overestimation. Beyond bias reduction, Q-function ensemble also facilitates exploration. Bootstrapped DQN (Osband et al., 2016) promoted behavioral diversity via bootstrap sampling and multi-headed Q-networks, while UCB Exploration (Chen et al., 2017) aggregated mean and variance from $K$ critics to guide action selection with an upper confidence bound. Similarly, UA-DDPG (Kanazawa et al., 2022) supported flexible ensemble sizes and encouraged exploration by prioritizing actions with high epistemic uncertainty. To further mitigate instability and balance exploration-exploitation, SUNRISE (Lee et al., 2021) was proposed as a unified ensemble RL framework, which re-weighted target Q-values based on ensemble uncertainty selects actions via UCB. Despite empirical success in standard RL settings, these methods are primarily designed for stationary, non-adversarial environments, leaving their application in robust adversarial RL largely unexplored. Under adversarial disturbances, Q-value estimates exhibit significant fluctuation, and simplistic pessimistic value estimation can hinder policy learning. Unlike prior work, we introduce a diversified critic ensemble to stabilize Q-learning against strategic perturbations and propose a time-varying decay uncertainty mechanism for Q-value aggregation, tailored to promote rapid exploration and policy learning in non-stationary adversarial RL training.
3 Problem Formulation
@@
We consider a robust adversarial RL framework modeled as a two-player $\gamma$ discounted zero-sum Markov game (Littman, 1994; Perolat et al., 2015): $\mathcal{M}=\langle\mathcal{S},\mathcal{A}_{p},\mathcal{A}_{a},\mathcal{P},\mathcal{R},\\ \gamma\rangle$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}_{p}$ and $\mathcal{A}_{a}$ are the action spaces of the protagonist and adversary, respectively, and $\mathcal{P}:\mathcal{S}\times\mathcal{A}_{p}\times\mathcal{A}_{a}\times\mathcal{S}\rightarrow\mathbb{R}$ defines the transition probability density, $\mathcal{R}:\mathcal{S}\times\mathcal{A}_{p}\times\mathcal{A}_{a}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward of both players, and $\gamma\in[0,1)$ is the discount factor. Our goal is to learn the policy of the protagonist (denoted by $\pi_{p}$) such that it can obtain maximum return in adversarial environments. Supposing the policy of the adversary is denoted by $\pi_{a}$, the optimization objective of the protagonist is to seek to maximize the following expected return:
@@
In the practical implementation, the adversarial training follows an alternating procedure: at each iteration, we first hold the protagonist’s policy fixed and update the adversary’s policy to minimize the expected return $J$, then conversely freeze the adversary’s policy and learn the protagonist’s policy to maximize the return. This procedure introduces training instability and convergence difficulties due to concurrent policy updates and adversarial perturbations. Our work specifically addresses these challenges through ensemble-based value estimation and adaptive uncertainty-aware policy optimization.
4 Methodology

### v2

@@ Abstract
Robust adversarial reinforcement learning has emerged as an effective paradigm for training agents to handle uncertain disturbance in real environments, with critical applications in sequential decision-making domains such as autonomous driving and robotic control. Within this paradigm, agent training is typically formulated as a zero-sum Markov game between a protagonist and an adversary to enhance policy robustness. However, the trainable nature of the adversary inevitably induces non-stationarity in the learning dynamics, leading to exacerbated training instability and convergence difficulties, particularly in high-dimensional complex environments. In this paper, we propose a novel approach, Uncertainty-Adaptive Critic Ensemble for robust adversarial Reinforcement learning (UACER), which consists of two components: 1) Diversified critic ensemble: A diverse set of $K$ critic networks is employed in parallel to stabilize Q-value estimation in robust adversarial reinforcement learning, reducing variance and enhancing robustness compared to conventional single-critic designs. 2) Time-varying Decay Uncertainty (TDU) mechanism: Moving beyond simple linear combinations, we propose a variance-derived Q-value aggregation strategy that explicitly incorporates epistemic uncertainty to adaptively regulate the exploration-exploitation trade-off while stabilizing the training process. Comprehensive experiments across several challenging MuJoCo control problems validate the superior effectiveness of UACER, outperforming state-of-the-art methods in terms of overall performance, stability, and efficiency.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
I Introduction
@@
Deep reinforcement learning (DRL) has established itself as a transformative paradigm for sequential decision-making, demonstrating remarkable success beyond gaming domains. Its applications now span real-world scenarios including precision robotic control [16, 8, 37, 33], adaptive healthcare systems [40, 32, 23], and autonomous vehicle navigation [38, 11, 46]. However, a critical gap persists: most DRL policies are trained in idealized simulated environments that fail to capture the complexity of physical world deployment. This issue becomes particularly pronounced in certain high-dimensional complex tasks such as robotic control. Even minor discrepancies in physical parameters, such as friction, mass distribution, or object appearance, can lead to significant performance deterioration when policies are deployed to real-world systems. More specifically, these policies frequently overfit to simulation-specific dynamics, resulting in catastrophic performance declines when confronted with inevitable real-world disturbances [14] or distribution shifts [5, 45]. Addressing this vulnerability, recent research [28, 43, 30] has prioritized the development of RL frameworks capable of counteracting adversarial disturbances and environmental variations.
@@
Robust adversarial reinforcement learning (RARL) [28, 31, 30] is a prominent class of methods in this research domain, specifically designed to enhance the stability and generalization capabilities of RL agents operating in uncertain or adversarially perturbed environments. The framework reformulates conventional RL as a two-player zero-sum Markov game [21, 27], comprising a protagonist agent that optimizes policy parameters to maximize cumulative returns and an adversary agent that systematically generates worst-case adversarial perturbations to minimize these returns. Whereas this minimax formulation theoretically enables agents to develop resilience against worst-case perturbations, its practical deployment poses significant methodological challenges [44]: (i) Non-Stationary Training Dynamics: The adversary’s evolving policy induces continuous perturbations, dynamically altering the protagonist’s observed state distribution and reward signals. This interaction creates a highly non-stationary learning environment where the protagonist’s optimization target shifts iteratively. As a result, policy updates exhibit high variance, frequently destabilizing the training process. (ii) Convergence Impediments: The cyclical adversarial training regime forces the protagonist to adapt to an ever-changing adversary policy before its own policy converges. Without a fixed optimization target, the protagonist remains trapped in a persistent reactive adaptation loop, leading to significantly prolonged convergence times due to continual policy adjustments, or even completed convergence failure in cases of extreme non-stationary conditions.
@@
To address the aforementioned stability and convergence problems, we propose Uncertainty-Adaptive Critic Ensemble robust adversarial Reinforcement learning (UACER), an effective general framework that advances conventional robust adversarial RL by systematically integrating critic ensemble methodology with uncertainty-adaptive policy optimization. This integration significantly improves both training stability and convergence speed in adversarial environments. The technical foundation of UACER addresses a fundamental limitation of single-critic architectures, which are particularly vulnerable to adversarial perturbations and distributional outliers, leading to unstable Q-value estimation [2, 26]. To overcome this limitation, UACER employs a parallel ensemble of critic networks, yielding three key benefits: Firstly, by aggregating Q-value estimates from multiple critics, the variance in Q-function estimation and policy gradient updates is reduced, thereby enhancing the stability of training; Secondly, the diversified value function approximation estimators effectively suppress training oscillations and improve robustness against adversarial disturbances; Thirdly, the resulting improvement in stability and robustness facilitates faster policy convergence during iterative adversarial interactions. For the final Q-value determination session, we present a Time-varying Decay Uncertainty (TDU) Q-value aggregation function to adaptively combine outputs from multiple critic networks for policy optimization. The TDU mechanism is designed to: (i) encourage extensive exploration during the initial training phases by maintaining a higher tolerance for optimistic uncertainty, and (ii) progressively transition toward more conservative value estimation in later stages to ensure stable convergence. This approach allows for a dynamically adaptive balance between exploration and exploitation, while also preserving resilience against adversarial perturbations throughout the learning process.
The contributions of this article are summarized as follows.
@@
    We propose a novel critic ensemble method for robust adversarial reinforcement learning, called UACER, which achieves more stable and accurate Q-value estimation by employing multiple base critic predictors.
@@
    We introduce the TDU function, a principled decision strategy for adaptively combining Q-value predictions from multiple critics for policy optimization. This mechanism dynamically balances exploration and exploitation while ensuring stable policy updates throughout the training process.
@@
    Quantitative experimental results on several complex MuJoCo tasks demonstrate that UACER consistently outperforms state-of-the-art methods, including QARL and related baselines, in terms of overall performance, stability, and efficiency.
II Related Work
II-A Robust Adversarial Reinforcement Learning
@@
The concept of adversarial training for robust RL was first established by RARL [28], which formulated the problem as a zero-sum Markov game between a protagonist agent and an adversarial opponent. This paradigm demonstrated that exposing agents to strategically generated perturbations during training could significantly enhance their robustness against real-world distributional shifts. Building upon this framework, subsequent research has developed specialized improvements targeting several key limitations. For instance, RARARL [25] incorporated explicit risk modeling to reduce catastrophic failures, while NR-MDP [34] expanded the adversary’s action space to generate more diverse and challenging disturbances. Further optimization advances introduced subsequently by MixedNE-LD [17] and Cen et al. [4] employed stochastic gradient Langevin dynamics and extragradient methods, respectively, to improve the robustness of adversarial training. More recently, A2P [22] proposed dynamically adjusting the intensity of action-space perturbations based on the agent’s current relative performance, enabling more adaptive adversarial training. Although these approaches have progressively addressed specific robustness challenges, they remain limited in handling the fundamental stability and convergence inherent in adversarial training. Our work advances this line of research by developing a unified framework that simultaneously improves training stability and convergence properties, while preserving and further enhancing the robustness benefits of adversarial training paradigms.
II-B Training Stability and Convergence Advances in RARL
@@
While RARL has markedly advanced the robustness of RL policies, its adversarial formulation introduces fundamental challenges in training stability and convergence. These limitations are not confined to complex settings but are inherent to the framework itself, as rigorously demonstrated by [44], who showed their persistence even in simple linear quadratic regulation problems. In response, recent research has primarily evolved along three directions: (i) Constrained optimization approaches, where DI-CARL [42] established stability guarantees through dissipation inequation constraint and $L_{2}$ gain performance in adversary RL, and Yu et al. [41] proposed an accelerated convergence method for linear quadratic games by employing Bregman divergence to effectively capture and adapt to the global structure of the constraint set. (ii) Curriculum learning methods enhance stability and robustness by progressively exposing the agent to escalating adversarial difficulty, notably through CAT [31], which employed a structured curriculum to systematically increase task complexity, thereby ensuring a stable and efficient training process. (iii) Hybrid frameworks such as QARL [30] that combine entropy regularization with quantal response equilibrium and use a temperature parameter to systematically manage adversary rationality in a curriculum learning manner, thereby accelerating convergence while easing the complexity of the saddle point optimization problem. In contrast, our work adopts an ensemble-based perspective that provides inherent stability to adversarial RL architecture through diversified Q-value estimation. This approach maintains compatibility with existing optimization strategies and further enhances the stability and convergence properties.
II-C Q-Function Ensemble in Reinforcement Learning
@@
The integration of multiple Q-value estimators, known as Q-function ensemble, has become a widely adopted approach to enhance the stability and robustness of value-based deep RL. Early ensemble methods in the DQN family primarily address overestimation bias through averaged or pessimistic selection strategies, such as the simple averaging in Averaged DQN [3] and the min-Q operator used in TD3 [12], Soft Actor-Critic (SAC) [15], and Maxmin Q-learning [19]. This conservative estimation approach was further extended by methods like MSG [13] and EDAC [1], which employed lower confidence bounds to systematically reduce overestimation. Beyond bias reduction, Q-function ensemble also facilitates exploration. Bootstrapped DQN [24] promoted behavioral diversity via bootstrap sampling and multi-headed Q-networks, while UCB Exploration [6] aggregated mean and variance from $K$ critics to guide action selection with an upper confidence bound. Similarly, UA-DDPG [18] supported flexible ensemble sizes and encouraged exploration by prioritizing actions with high epistemic uncertainty. To further mitigate instability and balance exploration-exploitation, SUNRISE [20] was proposed as a unified ensemble RL framework, which re-weighted target Q-values based on ensemble uncertainty and selected actions via UCB. Despite empirical success in standard RL settings, these methods are primarily designed for stationary, non-adversarial environments, leaving their application in robust adversarial RL largely unexplored. Under adversarial disturbances, Q-value estimates exhibit significant fluctuation, and simplistic pessimistic value estimation can hinder policy learning. Unlike prior work, we introduce a diversified critic ensemble to stabilize Q-learning against strategic perturbations and propose a time-varying decay uncertainty mechanism for Q-value aggregation, tailored to promote rapid exploration and policy learning in non-stationary adversarial RL training.
III Preliminaries and Notations
This section introduces the core terminology used in our paper, covering standard RL concepts, the basic RL framework SAC, and the RARL paradigm formulated as a two-player zero-sum game. The notations employed in this article are summarized in Table I.
III-A Standard RL and Soft Actor-Critic
Standard RL
RL addresses the problem of sequential decision-making and is typically formulated as a Markov Decision Process (MDP), in which an agent interacts with an environment over discrete time steps. At each time step $t$, the agent observes the current state $s_{t}$ and selects an action $a_{t}$ according to its policy $\pi$. In response, the environment provides a reward $r_{t}$ and transitions the agent to the subsequent state $s_{t+1}$. The objective of RL is to learn an optimal policy $\pi^{\ast}$ through repeated trial-and-error interactions and limited feedback, allowing the agent to effectively engage with the environment and maximize long-term cumulative rewards.
Soft Actor-Critic
Similar to the family of existing robust adversarial RL methods [28, 17, 30], we use SAC [15] as the underlying RL algorithm in this paper. SAC is an off-policy actor-critic method that extends the standard RL objective through entropy regularization, thereby enhancing exploration and policy robustness. The algorithm optimizes a stochastic policy that maximizes the expected discounted return together with an entropy-based regularization term. Formally, let $\pi_{\phi}(s_{t},a_{t})$ represent the policy network parameterized by $\phi$, and $Q_{\theta}(s_{t},a_{t})$ denote the Q-function network with parameters $\theta$. The Q-function is updated by minimizing the soft Bellman residual:
$\mathcal{L}_{Q}(\theta)=\mathbb{E}_{\tau_{t}\sim\mathcal{D}}\left[\big(Q_{\theta}(s_{t},a_{t})-R_{t}-\gamma V(s_{t+1})\big)^{2}\right],$ (1)
where $V(s_{t})=\mathbb{E}_{a_{t}\sim\pi_{\phi}}\big[Q_{\theta}(s_{t},a_{t})-\alpha\log\pi_{\phi}(a_{t}|s_{t})\big]$ is the soft state value function, $\tau_{t}$ is a transition, $\mathcal{D}$ is a replay buffer, and $\alpha$ is a temperature parameter. The policy is updated by minimizing the following objective:
$\mathcal{L}_{\pi}(\phi)=\mathbb{E}_{s_{t}\sim\mathcal{D}}\Big[\mathbb{E}_{a_{t}\sim\pi_{\phi}}\big[\alpha\log\pi(a_{t}|s_{t})-Q_{\theta}(s_{t},a_{t})\big]\Big].$ (2)
Collectively, these update rules establish a stable and sample-efficient learning foundation. In this work, we extend this framework to environments with uncertainty or adversarial disturbances through uncertainty-adaptive ensemble mechanisms.
TABLE I: Notations and their descriptions.
Notation Description
$\gamma$ discount factor
$t$ time step
$a_{p},a_{a}$ actions taken by the protagonist and adversary
$\theta_{k}^{p},\theta_{k}^{a}$ parameters of the $k^{\text{th}}$ critic networks for protagonist and adversary
$\phi_{p},\phi_{a}$ parameters of the actor networks for protagonist and adversary
$\mathcal{D}_{p},\mathcal{D}_{a}$ replay buffer for protagonist and adversary
$\pi_{p},\pi_{a}$ policies of the protagonist and adversary
$K$ number of ensemble critic networks
$\alpha$ temperature parameter in SAC
$\beta_{0}$ initial value of the time-decaying coefficient in TDU
$\beta_{\text{min}}$ a small constant for time-decaying coefficient in TDU
$\lambda$ decaying speed in TDU
NN total number of iterations
III-B RARL with Two-Player Zero-Sum Markov Games
@@
We consider an adversarial training framework for RL in adversarially perturbed environments, modeled as a two-player $\gamma$ discounted zero-sum Markov game [21, 27]: $\mathcal{M}=\langle\mathcal{S},\mathcal{A}_{p},\mathcal{A}_{a},\mathcal{P},\mathcal{R},\gamma\rangle$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}_{p}$ and $\mathcal{A}_{a}$ are the action spaces of the protagonist and adversary, respectively; $\mathcal{P}:\mathcal{S}\times\mathcal{A}_{p}\times\mathcal{A}_{a}\times\mathcal{S}\rightarrow\mathbb{R}$ defines the transition probability density, and $\mathcal{R}:\mathcal{S}\times\mathcal{A}_{p}\times\mathcal{A}_{a}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward of both players. Our goal is to learn the policy of the protagonist (denoted by $\pi_{p}$) such that it can obtain maximum $\gamma$ discounted return in adversarial environments. Supposing the policy of the adversary is denoted by $\pi_{a}$, the optimization objective of the protagonist is to seek to maximize the following expected return:
@@
In the practical implementation, the adversarial training follows an alternating procedure: at each iteration, we first hold the protagonist’s policy fixed and update the adversary’s policy to minimize the expected return $J$, then conversely freeze the adversary’s policy and learn the protagonist’s policy to maximize the return. This procedure introduces training instability and convergence difficulties due to concurrent policy updates and adversarial perturbations. Our work specifically addresses these challenges through ensemble-based value estimation and uncertainty-adaptive policy optimization.
IV Methodology
Figure 1: Schematic of the proposed UACER method, comprising: (a) robust adversarial RL components for protagonist-adversary interaction; (b) uncertainty-aware critic ensemble training architecture with $K$ parallel critics and TDU Q-value estimation; (c) detailed decomposition of TDU function.